### v1

# MusiCRS: Benchmarking Audio-Centric Conversational Recommendation

Abstract
Conversational recommendation has advanced rapidly with large language models (LLMs), yet music remains a uniquely challenging domain where effective recommendations require reasoning over audio content beyond what text or metadata can capture. We present MusiCRS, the first benchmark for audio-centric conversational recommendation that links authentic user conversations from Reddit with corresponding audio tracks. MusiCRS contains 477 high-quality conversations spanning diverse genres (classical, hip-hop, electronic, metal, pop, indie, jazz) with 3,589 unique musical entities and audio grounding via YouTube links. MusiCRS enables evaluation across three input modality configurations: audio-only, query-only, and audio+query (multimodal), allowing systematic comparison of audio-LLMs, retrieval models, and traditional approaches. Our experiments reveal that current systems rely heavily on textual signals and struggle with nuanced audio reasoning. This exposes fundamental limitations in cross-modal knowledge integration where models excel at dialogue semantics but cannot effectively ground abstract musical concepts in actual audio content. To facilitate progress, we release the MusiCRS dataset [^1] https://huggingface.co/datasets/rohan2810/MusiCRS, evaluation code [^2] https://github.com/rohan2810/musiCRS, and comprehensive baselines.
Index Terms—  Conversational Recommendation, Music Recommendation, Music Information Retrieval, Multimodal Learning
1 Introduction
Dataset Source Type Real Audio Ground. Rec. Multi.
CPCD (Playlist) [1] Crowdsourced
TalkTheWalk [2] Synthetic (LLM+lists)
LP-MusicDialog [3] Synthetic (LLM-gen)
Audio Dialogues [4] Synthetic (captions)
LastFM [5] User logs
MusiCRS In-the-wild (Reddit)
Fig. 1: Limitations of existing conversational recommendation approaches (top) and comparison of music recommendation datasets (bottom). MusiCRS is the only benchmark combining authentic conversations, audio grounding, ground truth annotations, recommendation evaluation, and multimodal capabilities.
Conversational recommender systems have emerged as an important testbed for large language models (LLMs), enabling multi-turn interactions that track user intent and generate contextually relevant recommendations. While substantial progress has been made in domains like movies and e-commerce [6, 7, 8], music recommendation presents unique challenges. Listener preferences are influenced not only by dialogue context but also by musical characteristics such as rhythm, timbre, instrumentation, and production style. These subtle musical dimensions are difficult to capture through text or metadata alone [9], making music a demanding setting for studying multimodal conversational recommendation. Current benchmarks predominantly focus on movies or general domains [10], while music datasets emphasize metadata, tags, or listening logs [11] rather than conversations referencing actual recordings. This gap prevents evaluation of whether systems can bridge natural conversational cues with the musical characteristics driving human preference, limiting progress in multimodal music understanding.
Fig. 2: Representative examples from MusiCRS showing Reddit conversations, derived queries, audio, and candidates across genres.
To address this gap, we introduce MusiCRS, the first benchmark for audio-centric conversational recommendation that explicitly links conversations with corresponding audio tracks. Built from large-scale Reddit discussions where users naturally exchange music suggestions, the dataset grounds each recommendation query in a referenced track to reflect authentic listening contexts. As illustrated in Fig. 1, MusiCRS integrates conversational signals, audio grounding, and recommendation context within a single benchmark, enabling the evaluation of whether models can jointly interpret dialogue and musical content rather than relying on metadata or popularity cues.
MusiCRS contains 477 high-quality Reddit discussions spanning seven major genres, with each conversation grounded in actual audio through validated YouTube links. This design captures authentic music discovery scenarios and enables rigorous evaluation of cross-modal understanding in conversational recommendation.
MusiCRS enables evaluation across three input modality configurations that systematically assess the contribution of different information sources. The audio-only configuration tests models using only the referenced audio tracks, isolating the impact of musical content. The query-only configuration evaluates models relying solely on conversational text, aligned with traditional text-based recommendation systems. The audio+query configuration combines both modalities, testing recent advances in audio-LLMs [12, 13, 14] and multimodal retrieval models [15, 16]. This design enables systematic analysis of how different information sources contribute to music recommendation quality across generative, retrieval-based, and traditional approaches. Our contributions are threefold:
  • We introduce MusiCRS, the first benchmark for audio-centric conversational music recommendation, built from 477 Reddit discussions across seven genres with comprehensive audio grounding via validated YouTube links.
  • We establish rigorous evaluation protocols across three input modalities and provide comprehensive benchmarking of audio-LLMs, retrieval models, and traditional approaches, revealing substantial performance gaps in current systems.
  • We uncover a fundamental limitation in current multimodal models: optimal performance frequently occurs in single-modality settings, revealing that models fail to effectively ground abstract musical concepts in audio content, exposing critical challenges in cross-modal knowledge integration.
2 MusiCRS Dataset
2.1 Dataset Construction
Large-Scale Reddit Mining and Filtering. We collected 2.7 million Reddit submissions and 28.5 million comments from high-activity, music-oriented subreddits including r/classicalmusic, r/hiphopheads, r/electronicmusic, r/metal, r/popheads, r/indieheads, and r/jazz, following prior large-scale Reddit-based efforts in conversational AI [17, 18]. We applied a three-stage filtering pipeline to ensure both musical grounding and conversational quality. First, only posts with valid YouTube links were retained, yielding $46,218$ threads linked to external music resources. Second, threads with fewer than three first-level replies or comments under five characters were removed, leaving $10,167$ multi-turn discussions with substantive context [10]. Finally, we discarded low-interaction threads during human annotation, ensuring the final dataset contained coherent and contextually rich exchanges.
Music Entity Extraction and Query Generation. We use LLMs to transform Reddit discussions into structured inputs for downstream tasks. Entity extraction was performed with Qwen2.5-7B [19], which identified and categorized songs, artists, and albums from post titles, bodies, and top-level comments. In parallel, gemma-3-12b-it [20] generated concise, context-preserving queries from each post, capturing user intent to support recommendation and retrieval.
Query Normalization and Enhanced Candidate Selection. For evaluation, we construct candidate pools that balance ground truth inclusion with realistic candidates: each conversation’s candidate pool combines the top $10$ ground truth entities (ranked by comment upvotes) with approximately $90$ additional entities randomly sampled from the same subreddit, then shuffled to eliminate positional bias. This approach ensures comprehensive evaluation coverage while maintaining topical coherence within music communities.
Corpus Refinement and Audio Grounding. We refine the dataset by retaining conversations with 2–30 musical entities, focusing on substantive discussions while eliminating noise. Comments were limited to 5 entities to prevent verbosity and topic drift [21]. Ground truth entities are extracted from comments and ranked by upvote scores. By incorporating popularity signals while maintaining diverse subreddit representation, this approach balances mainstream and niche musical content [22]. YouTube links were validated and processed into audio clips, grounding each conversation in authentic music content for evaluation. Fig. 2 illustrates genre-specific conversations, including the derived query, audio clip, and candidate pools.
Human Validation. To ensure conversational and musical grounding, four graduate students manually validated $1,623$ conversations. Each was categorized as music, pop culture, or other, with YouTube link verification and entity relevance review. During this process, $41.3\%$ of URLs were discarded as invalid while $10.5\%$ were corrected or added through manual verification. Entity refinement removed $20.0\%$ of irrelevant entities and recovered $0.3\%$ missing ones, improving overall precision. Additionally, $31.1\%$ of conversations were reassigned to correct categories, ensuring proper alignment with the music-focused benchmark. Spam and off-topic content was removed throughout this validation process. For the final MusiCRS benchmark, we retained only the conversations classified as music, resulting in $477$ high-quality music-focused conversational threads.
Fig. 3: Dataset genre and entity coverage (left) and dialogue structure statistics (right). MusiCRS encompasses diverse genres, artists, and song distributions, and a wide range of conversation styles. Macro-genre abbreviations: $M$ (Modern & Mainstream: hiphop, pop); IM (Instrumental / Art Music: classical, jazz); AE (Alternative / Experimental: indie, metal, electronic).
2.2 Dataset Statistics
The curated MusiCRS benchmark contains 477 conversational threads spanning 3,589 musical entities, with 8 ground truth entities per conversation on average. Each conversation includes 100 candidates for evaluation, selected as described above by combining ground-truth items with subreddit-specific distractors to create realistic recommendation lists. Figure 3 summarizes genre balance and subreddit proportions: classical music (25.8%), hip-hop (21.0%), indie (12.8%), electronic (11.7%), metal (10.3%), pop (9.4%), and jazz (9.0%). This distribution ensures comprehensive coverage across mainstream and niche musical content, capturing variation in conversational styles, knowledge depth, and recommendation contexts, establishing MusiCRS as a robust benchmark for conversational, multimodal, and content-based music recommendation.
3 Experiments
To evaluate the effectiveness of audio-centric conversational recommendation systems on the MusiCRS dataset, we conduct comprehensive experiments across multiple model paradigms and input modalities. Our analysis focuses on three key dimensions: (1) comparing generative, retrieval-based, and traditional recommendation methods; (2) evaluating the impact of different input modalities (audio, text query, and multimodal combinations); and (3) assessing model performance across diverse musical genres and subreddits.
3.1 Experimental Setup
We frame experiments as ranking tasks across three modalities: audio-only, query-only, and audio+query. Each query uses a 300-second audio budget across up to 10 clips (30 seconds for SALMONN). Audio is resampled to model-appropriate rates (16kHz for SALMONN, 48kHz for CLAP). We employ Recall, nDCG, and MRR as primary metrics, aggregated across the dataset and per-subreddit. Generative models use default generative settings with increased max new tokens (typically 512 tokens) to accommodate full candidate rankings. Retrieval models compute cosine similarity between query embeddings and candidate song embeddings, with late fusion for multimodal scenarios. We evaluate seven generative models (Qwen2-Audio-7B [12], Qwen2.5-Omni-7B [23], SALMONN-7B/13B [13], FUTGA [24], Audio Flamingo 3 [14], Phi-4-Multimodal [25]), two retrieval systems (CLAP [26], CoLLAP [16]), and two traditional baselines (subreddit-specific popularity, neighborhood-based collaborative recommendation).
Model Modality Classical Hip-Hop Pop Electronic Metal Indie Jazz Overall
[…]
Qwen2-Audio Audio 15.83/12.10 14.20/11.40 17.47/13.20 17.90/11.75 18.01/11.02 16.74/12.48 19.19/10.98 16.53/11.86
Query 16.15/10.97 17.16/13.01 18.02/13.33 10.70/7.98 13.73/10.05 17.18/13.86 16.00/10.72 15.77/11.52
Combined 16.38/12.71 18.39/14.03 15.32/12.29 14.58/10.23 17.40/13.26 15.61/12.63 21.46/13.76 16.95/12.80
Qwen2.5-Omni Audio 14.80/10.79 19.24/13.28 16.47/12.07 19.79/15.22 22.47/13.57 21.08/16.62 28.09/16.36 19.26/13.48
Query 14.07/12.04 18.91/14.10 23.38/17.59 18.00/13.50 16.45/12.62 20.39/14.49 22.48/16.73 18.24/13.96
Combined 26.53/19.33 17.74/13.55 18.92/13.14 25.29/19.03 23.61/16.91 17.30/14.23 22.00/15.02 21.93/16.21
SALMONN-7B Audio 17.88/12.89 25.12/16.85 15.41/10.66 16.79/12.16 21.41/15.08 22.24/17.07 20.76/14.35 20.22/14.31
Query 19.17/13.01 19.33/14.05 14.07/9.68 23.55/13.66 17.83/12.60 19.22/13.55 13.49/8.63 18.60/12.62
Combined 21.70/15.06 19.91/13.15 15.74/11.71 18.31/12.97 16.03/10.60 18.99/15.75 23.35/15.06 19.58/13.73
SALMONN-13B Audio 17.12/12.08 22.31/15.14 16.39/11.13 18.12/12.31 19.15/14.57 19.94/14.98 25.18/16.26 19.55/13.66
Query 17.96/12.62 19.78/13.90 15.38/10.63 22.32/13.08 19.81/14.34 21.91/15.99 17.80/10.95 19.29/13.21
Combined 18.48/12.86 19.23/12.57 16.41/11.29 17.48/12.50 16.98/10.64 18.70/15.57 25.47/17.39 18.83/13.13
FUTGA Audio 18.89/11.72 17.84/12.25 19.25/13.69 17.60/12.72 19.79/12.53 22.60/14.98 20.32/14.07 19.25/12.84
Query 19.11/12.08 18.42/11.85 22.65/14.82 14.97/9.79 16.96/10.82 20.77/14.23 20.73/10.40 18.95/12.02
Combined 19.58/11.89 17.56/11.61 18.89/12.87 19.25/12.29 15.36/10.51 20.48/13.96 18.88/11.24 18.67/12.04
Audio Flamingo 3 Audio 17.30/12.38 14.20/10.78 13.61/10.33 16.02/11.79 22.41/12.35 21.62/15.54 19.48/12.88 17.42/12.23
Query 17.02/13.52 17.77/12.49 17.61/12.62 17.59/12.93 19.33/14.71 20.77/14.38 20.72/12.91 18.35/13.33
Combined 16.38/12.93 17.56/12.12 17.01/12.69 14.82/10.73 16.70/10.33 20.97/15.00 25.52/14.93 17.95/12.66
Phi-4-Multimodal Audio 21.86/15.08 20.72/14.57 14.56/10.01 18.17/13.53 18.75/11.38 22.95/16.53 18.72/10.68 20.04/13.72
Query 24.38/18.57 17.01/13.37 16.69/11.45 12.82/8.58 15.82/10.70 29.85/17.40 17.31/10.52 19.93/13.95
Combined 19.94/12.96 16.91/11.63 17.91/12.59 18.23/12.59 15.80/9.79 20.68/14.80 22.31/15.72 18.79/12.76
[…]
CLAP Audio 22.38/15.68 21.22/15.14 18.86/14.93 15.12/12.11 26.42/16.25 20.56/15.41 22.51/13.40 21.15/14.90
Query 23.56/16.87 23.45/16.67 18.50/13.78 18.23/12.82 23.74/15.78 23.43/17.09 26.56/15.98 22.71/15.90
Combined 23.79/16.96 22.10/15.85 18.52/15.24 17.26/12.19 25.00/16.43 22.14/17.18 27.61/15.23 22.43/15.82
CoLLAP Audio 17.23/12.22 19.90/14.49 19.75/13.52 19.04/12.75 20.24/14.11 18.17/14.13 27.75/15.52 19.62/13.62
Query 19.26/12.70 20.30/14.75 20.38/15.96 22.11/13.87 20.19/12.07 21.65/16.44 25.14/14.35 20.85/14.14
Combined 18.68/12.86 21.09/15.08 21.88/15.98 21.17/14.63 22.35/13.04 18.42/12.66 23.08/14.02 20.52/13.92
[…]
Popularity Query 16.41/11.21 15.15/10.18 16.52/12.30 16.73/10.10 14.92/9.62 15.05/11.67 23.59/13.70 16.51/11.09
Neighborhood Query 20.12/13.53 17.47/10.92 15.85/10.35 13.76/7.92 11.91/6.67 8.96/5.59 4.30/2.38 14.72/9.30
Table 1: Performance comparison across music genres and input modalities. Values represent Recall@20/nDCG@20. Bold values indicate the best-performing modality configuration for each model, and underlined values show the top-performing model for each genre.
3.2 Research Questions and Analysis
Our comprehensive evaluation addresses three key research questions that illuminate fundamental challenges in multimodal music recommendation (Table 1):
RQ1: How do different input modalities impact music recommendation performance?
Our analysis reveals that multimodal approaches do not consistently outperform single-modality configurations. CLAP achieves optimal performance with query-only input (22.71% Recall@20) compared to audio-only (21.15%), while Qwen2.5-Omni demonstrates superior performance with audio+query combination (21.93% vs. 19.26% audio-only and 18.24% query-only). However, only 3 of 9 evaluated models achieve their best performance in multimodal settings. This finding aligns with recent work [27, 28] demonstrating that current multimodal models exhibit strong reliance on textual signals due to architectural limitations in cross-modal integration, suggesting that existing fusion mechanisms inadequately exploit complementary information across modalities.
Case Study(RQ1) Q: What songs evoke overwhelming sorrow and emptiness, similar to Mozart’s Lacrimosa or the opening of Elfen Lied?
Ground Truth Candidate(s):
Nymphs in the woods, Lamentations of Jeremiah, Ne irascaris & Civitas, The One, Elfen Lied, Lost in Paradise, Speak to Me, Adagio, Miserere, 2nd movement Andante
Query (Top-5):
Wake Up, Numb, My Immortal, One More Light, Fix You
Audio (Top-5):
Ave Maria, Clair de Lune, Moonlight Sonata, Canon in D, Adagio for Strings
Combined (Top-5):
Lamentations of Jeremiah, Daphnis et Chloé, Elfen Lied, Requiem, Miserere
RQ2: How do different model paradigms compare in music recommendation?
Retrieval-based approaches demonstrate superior performance compared to generative models across all metrics. CLAP achieves the highest overall performance (22.71% Recall@20), followed by CoLLAP (20.85%), while the best-performing generative model (Qwen2.5-Omni) reaches 21.93%. Traditional methods exhibit substantially lower performance, with popularity-based recommendation achieving 16.51% and neighborhood-based filtering 14.72%, highlighting their inadequacy for audio-centric recommendation tasks that require understanding musical content beyond popularity signals. Within generative models, performance varies considerably (17.42% to 21.93%), whereas retrieval models maintain more consistent performance across input modality configurations. These results indicate that learned embedding representations provide more robust recommendation signals than generative reasoning for music recommendation tasks.
Case Study (RQ2) Q: Following a recommendation from a Benjamin Zander TED Talk, what classical works would you suggest for a newcomer seeking emotionally evocative pieces?
Ground Truth Candidates: Adagio for Strings, Symphony No.6, 7th Symphony (Second Movement), Moonlight Sonata, Op.18, Six Etudes-tableaux, Adagio-Lacrimosa, Elgar Cello Concerto, Nocturne Op. 55 No.1 in F minor, Ballade in G Minor
Traditional (Popularity) (Top-5):
Fantasy in F minor, Rosamunde Quartetsatz,
Nocturne Op. 55 No. 1 in F minor, Late-Romantic, Lux Aurumque
Retrieval (CLAP) (Top-5):
Waldstein Sonata, Shostakovich Concerto,
Six Etudes-tableaux, 7th Symphony (Second Movement), Piano Concerto No. 2 Mvt. 3
Generative (Qwen2.5-Omni) (Top-5):
Adagio for Strings, Rach 1 Ossia, Ballade in G Minor, 7th Symphony (Second Movement), La Mer
Fig. 4: Mean Reciprocal Rank (MRR) comparison across genres and model types.
RQ3: What genre-specific patterns emerge?
Performance analysis (Figure 4) reveals variation that reflects both domain-specific challenges and model capabilities. Jazz consistently achieves the highest performance across most models (28.09% for Qwen2.5-Omni), while Classical and Hip-Hop genres present persistent challenges across all approaches. Metal demonstrates strong performance with retrieval models (CLAP: 26.42% audio-only) but exhibits weaker results with generative approaches. Electronic and Indie genres show favorable performance with generative models, with several achieving >20% Recall@20. These patterns suggest that genre-specific audio characteristics and linguistic conventions differentially impact model performance, with Jazz’s distinctive musical features being more readily captured by current architectures.
4 Conclusion
We introduced MusiCRS, the first benchmark explicitly designed for audio-centric conversational recommendation, grounded in both conversational context and corresponding audio. Through comprehensive evaluation across diverse recommendation paradigms and genres, our experiments highlighted critical challenges faced by current multimodal and conversational models, particularly in nuanced audio-contextual reasoning. We release MusiCRS to advance research in multimodal music recommendation.