### v2

# MusiCRS: Benchmarking Audio-Centric Conversational Recommendation

Abstract
Conversational recommendation has advanced rapidly with large language models (LLMs), yet music remains a uniquely challenging domain in which effective recommendations require reasoning over audio content beyond what text or metadata can capture. We present MusiCRS, the first benchmark for audio-centric conversational recommendation that links authentic user conversations from Reddit with corresponding tracks. MusiCRS includes 477 high-quality conversations spanning diverse genres (classical, hip-hop, electronic, metal, pop, indie, jazz), with 3,589 unique musical entities and audio grounding via YouTube links. MusiCRS supports evaluation under three input modality configurations: audio-only, query-only, and audio+query, allowing systematic comparison of audio-LLMs, retrieval models, and traditional approaches. Our experiments reveal that current systems struggle with cross-modal integration, with optimal performance frequently occurring in single-modality settings rather than multimodal configurations. This highlights fundamental limitations in cross-modal knowledge integration, as models excel at dialogue semantics but struggle when grounding abstract musical concepts in audio. To facilitate progress, we release the MusiCRS dataset [^1] https://huggingface.co/datasets/rohan2810/MusiCRS, evaluation code [^2] https://github.com/rohan2810/musiCRS, and comprehensive baselines.
Index Terms—  Conversational Recommendation, Music Recommendation, Music Information Retrieval, Multimodal Learning
1 Introduction
Dataset Source Type Real Audio Ground. Rec. Multi.
CPCD (Playlist) [1] Crowdsourced
TalkTheWalk [2] Synthetic (LLM+lists)
LP-MusicDialog [3] Synthetic (LLM-gen)
Audio Dialogues [4] Synthetic (captions)
LastFM [5] User logs
MusiCRS In-the-wild (Reddit)
Fig. 1: Limitations of existing conversational recommendation approaches (top) and comparison of music recommendation datasets (bottom). MusiCRS is the only benchmark combining authentic conversations, audio grounding, ground truth annotations, recommendation evaluation, and multimodal capabilities.
Traditional music recommender systems rely on collaborative filtering and metadata but ignore the conversational dynamics through which users naturally express preferences. Conversational recommender systems powered by large language models (LLMs) address this limitation, enabling multi-turn interactions that track user intent and produce contextually relevant recommendations. While substantial progress has been made in domains like movies and e-commerce [6, 7, 8], music recommendation presents unique challenges. Listener preferences depend not only on dialogue context but also on musical characteristics such as rhythm, timbre, instrumentation, and production style. These subtle musical dimensions are difficult to capture through text or metadata alone [9], making music an ideal testbed for multimodal conversational recommendation. Current benchmarks predominantly focus on movies or general domains [10], while music datasets emphasize metadata, tags, or listening logs [5] rather than conversations referencing actual recordings. This gap prevents the evaluation of whether systems can bridge natural conversational cues with the musical characteristics driving human preference, limiting progress in multimodal music understanding.
Fig. 2: Representative examples from MusiCRS showing Reddit conversations, derived queries, audio, and candidates across genres.
To address this gap, we introduce MusiCRS, the first benchmark for audio-centric conversational recommendation that links conversations with corresponding audio tracks. Built from large-scale Reddit discussions where users naturally exchange music suggestions, the dataset grounds each query in a referenced track, reflecting real-world music discovery. As illustrated in Fig. 1, existing approaches face complementary limitations: traditional systems ignore conversational dynamics, while text-only conversational systems miss real audio grounding. MusiCRS addresses both by integrating dialogue context with authentic audio within a single benchmark. MusiCRS contains 477 high-quality Reddit discussions spanning seven genres, with each conversation grounded in audio through validated YouTube links. This design enables rigorous evaluation of cross-modal understanding in conversational music recommendation.
MusiCRS enables evaluation across three input modality configurations that systematically assess the contribution of different information sources. The audio-only configuration tests models using only the referenced audio tracks, isolating the impact of musical content. The query-only configuration evaluates models relying solely on conversational text, aligned with traditional text-based recommendation systems. The audio+query configuration combines both modalities, testing recent advances in audio-LLMs [11, 12, 13] and multimodal retrieval models [14, 15]. This design enables systematic analysis of how different information sources contribute to music recommendation quality across generative, retrieval-based, and traditional approaches. Our contributions are threefold:
  • We introduce MusiCRS, the first benchmark for audio-centric conversational music recommendation, built from 477 Reddit discussions across seven genres with comprehensive audio grounding via validated YouTube links.
  • We establish rigorous evaluation protocols across three input modalities and provide comprehensive benchmarking of audio-LLMs, retrieval models, and traditional approaches, revealing substantial performance gaps in current systems.
  • We uncover a fundamental limitation in current multimodal models: optimal performance frequently occurs in single-modality settings, showing that models fail to ground abstract musical concepts in audio content.
2 MusiCRS Dataset
2.1 Dataset Construction
Large-Scale Reddit Mining and Filtering. We collected 2.7 million Reddit submissions and 28.5 million comments from high-activity, music-oriented subreddits, including r/classicalmusic, r/hiphopheads, r/electronicmusic, r/metal, r/popheads, r/indieheads, and r/jazz, following prior large-scale Reddit-based efforts in conversational AI [16, 17]. We applied a three-stage filtering pipeline to ensure both musical grounding and conversational quality. First, only posts with valid YouTube links were retained, yielding $46,218$ threads linked to external music resources. Second, threads with fewer than three first-level replies or comments under five characters were removed, leaving $10,167$ multi-turn discussions with substantive context [10]. Finally, we discarded low-interaction threads during human annotation, ensuring the final dataset contained coherent and contextually rich exchanges.
Music Entity Extraction and Query Generation. We used LLMs to transform Reddit discussions into structured inputs for downstream tasks. Entity extraction was performed with Qwen2.5-7B [18], identifying and categorizing songs, artists, and albums from post titles, bodies, and top-level comments. In parallel, gemma-3-12b-it [19] generated concise, context-preserving queries from each post, capturing user intent to support recommendation and retrieval.
Query Normalization and Enhanced Candidate Selection. For evaluation, we constructed candidate pools that include the ground truth entities while remaining realistic. For each conversation, we take the top $10$ ground-truth entities (ranked by comment upvotes) and add about $90$ additional entities randomly sampled from the same subreddit. We then shuffle the pool to remove positional bias. This design provides broad evaluation coverage while preserving topical coherence within each music community.
Corpus Refinement and Audio Grounding. We refine the dataset by retaining conversations with 223030 musical entities, focusing on substantive discussions while eliminating noise. We cap each comment at $5$ entities to prevent verbosity and topic drift [20]. Ground truth entities are extracted from comments and ranked by upvote score. We use upvotes as a popularity signal while maintaining diverse subreddit representation, balancing mainstream and niche musical content [21]. YouTube links were validated and segmented into audio clips, grounding each conversation in authentic music content for evaluation. Fig. 2 illustrates genre-specific conversations, including the derived query, audio clip, and candidate pools.
Human Validation. To ensure conversational quality and audio grounding, four graduate student annotators manually validated $1,623$ conversations. Each was categorized as music, pop culture, or other, including YouTube link verification and entity relevance review. During this process, $41.3\%$ of URLs were discarded as invalid, while $10.5\%$ were corrected or added through manual verification. Entity refinement removed $20.0\%$ of irrelevant entities and recovered $0.3\%$ of missing entities, improving overall precision. Additionally, $31.1\%$ of conversations were reassigned to the correct categories, ensuring proper alignment with the music-focused benchmark. Spam and off-topic content were removed throughout this validation process. For the final MusiCRS benchmark, we retained only the conversations classified as music, resulting in $477$ high-quality music-focused conversational threads.
Fig. 3: Dataset genre and entity coverage (left) and dialogue structure statistics (right). MusiCRS encompasses diverse genres, artists, and song distributions, and a wide range of conversation styles. Macro-genre abbreviations: $M$ (Modern & Mainstream: hip-hop, pop); IM (Instrumental / Art Music: classical, jazz); AE (Alternative / Experimental: indie, metal, electronic).
2.2 Dataset Statistics
The curated MusiCRS benchmark contains $477$ conversational threads spanning $3,589$ musical entities, with an average of $8$ ground truth entities per thread. Each thread includes $100$ evaluation candidates, constructed by combining the ground-truth items with subreddit-specific distractors to form realistic recommendation lists. Fig. 3 summarizes genre balance and subreddit proportions: Classical (25.8%), Hip-Hop (21.0%), Indie (12.8%), Electronic (11.7%), Metal (10.3%), Pop (9.4%), and Jazz (9.0%). This distribution ensures comprehensive coverage across mainstream and niche musical content, capturing variation in conversational styles, knowledge depth, and recommendation contexts, establishing MusiCRS as a robust benchmark for conversational, multimodal, and content-based music recommendation.
Case Study (RQ1) Q: What songs evoke overwhelming sorrow and emptiness, similar to Mozart’s Lacrimosa or the opening of Elfen Lied? Ground Truth: Nymphs in the woods, Lamentations of Jeremiah, Ne irascaris & Civitas, The One, Elfen Lied, Lost in Paradise, Speak to Me, Adagio, Miserere, 2nd movement Andante Query (Top-5):
Wake Up, Numb, My Immortal, One More Light, Fix You
Audio (Top-5):
Ave Maria, Clair de Lune, Moonlight Sonata, Canon in D, Adagio for Strings
Combined (Top-5):
Lamentations of Jeremiah, Daphnis et Chloé, Elfen Lied, Requiem, Miserere
3 Experiments
To evaluate the effectiveness of audio-centric conversational recommendation systems on the MusiCRS dataset, we conduct comprehensive experiments across multiple model paradigms and input modalities. Our analysis focuses on three key dimensions: (1) comparing generative, retrieval-based, and traditional recommendation methods; (2) evaluating the impact of different input modalities (audio, text query, and multimodal combinations); and (3) assessing model performance across diverse musical genres and subreddits.
Case Study (RQ2) Q: Following a recommendation from a Benjamin Zander TED Talk, what classical works would you suggest for a newcomer seeking emotionally evocative pieces? Ground Truth: Adagio for Strings, Symphony No.6, 7th Symphony (Second Movement), Moonlight Sonata, Op.18, Six Etudes-tableaux, Adagio-Lacrimosa, Elgar Cello Concerto, Nocturne Op. 55 No.1 in F minor, Ballade in G Minor Traditional (Popularity) (Top-5):
Fantasy in F minor, Rosamunde Quartetsatz, Nocturne Op. 55 No. 1 in F minor, Late-Romantic, Lux Aurumque
Retrieval (CLAP) (Top-5):
Waldstein Sonata, Shostakovich Concerto, Six Etudes-tableaux, 7th Symphony (Second Movement),Piano Concerto No.2 Mvt.3
Generative (Qwen2.5-Omni) (Top-5):
Adagio for Strings, Rach 1 Ossia, Ballade in G Minor, 7th Symphony (Second Movement), La Mer
3.1 Experimental Setup
We frame experiments as ranking tasks across three modalities: audio-only, query-only, and audio+query. Each query uses a $300$-second audio budget across up to $10$ clips ($30$ seconds for SALMONN). Audio is resampled to model-appropriate rates (1616kHz for SALMONN, 4848kHz for CLAP). We employ Recall, nDCG, and MRR as primary metrics, aggregated across the dataset and per-subreddit. Generative models use default generative settings with increased max_new_tokens (typically $512$ tokens) to accommodate full candidate rankings. Retrieval models compute cosine similarity between query embeddings and candidate song embeddings, with late fusion for multimodal scenarios. We evaluate seven generative models (Qwen2-Audio-7B [11], Qwen2.5-Omni-7B [22], SALMONN-7B/13B [12], FUTGA [23], Audio Flamingo 3 [13], Phi-4-Multimodal [24]), two retrieval systems (CLAP [25], CoLLAP [15]), and two traditional baselines (subreddit-specific popularity, neighborhood-based recommendation).
Model Modality Classical Hip-Hop Pop Electronic Metal Indie Jazz Overall
Generative Models
Qwen2-Audio-7B Audio 15.83/12.10 14.20/11.40 17.47/13.20 17.90/11.75 18.01/11.02 16.74/12.48 19.19/10.98 16.53/11.86
Query 16.15/10.97 17.16/13.01 18.02/13.33 10.70/07.98 13.73/10.05 17.18/13.86 16.00/10.72 15.77/11.52
Combined 16.38/12.71 18.39/14.03 15.32/12.29 14.58/10.23 17.40/13.26 15.61/12.63 21.46/13.76 16.95/12.80
Qwen2.5-Omni-7B Audio 14.80/10.79 19.24/13.28 16.47/12.07 19.79/15.22 22.47/13.57 21.08/16.62 28.09/16.36 19.26/13.48
Query 14.07/12.04 18.91/14.10 23.38/17.59 18.00/13.50 16.45/12.62 20.39/14.49 22.48/16.73 18.24/13.96
Combined 26.53/19.33 17.74/13.55 18.92/13.14 25.29/19.03 23.61/16.91 17.30/14.23 22.00/15.02 21.93/16.21
SALMONN-7B Audio 17.88/12.89 25.12/16.85 15.41/10.66 16.79/12.16 21.41/15.08 22.24/17.07 20.76/14.35 20.22/14.31
Query 19.17/13.01 19.33/14.05 14.07/9.68 23.55/13.66 17.83/12.60 19.22/13.55 13.49/8.63 18.60/12.62
Combined 21.70/15.06 19.91/13.15 15.74/11.71 18.31/12.97 16.03/10.60 18.99/15.75 23.35/15.06 19.58/13.73
SALMONN-13B Audio 17.12/12.08 22.31/15.14 16.39/11.13 18.12/12.31 19.15/14.57 19.94/14.98 25.18/16.26 19.55/13.66
Query 17.96/12.62 19.78/13.90 15.38/10.63 22.32/13.08 19.81/14.34 21.91/15.99 17.80/10.95 19.29/13.21
Combined 18.48/12.86 19.23/12.57 16.41/11.29 17.48/12.50 16.98/10.64 18.70/15.57 25.47/17.39 18.83/13.13
FUTGA Audio 18.89/11.72 17.84/12.25 19.25/13.69 17.60/12.72 19.79/12.53 22.60/14.98 20.32/14.07 19.25/12.84
Query 19.11/12.08 18.42/11.85 22.65/14.82 14.97/09.79 16.96/10.82 20.77/14.23 20.73/10.40 18.95/12.02
Combined 19.58/11.89 17.56/11.61 18.89/12.87 19.25/12.29 15.36/10.51 20.48/13.96 18.88/11.24 18.67/12.04
Audio Flamingo 3 Audio 17.30/12.38 14.20/10.78 13.61/10.33 16.02/11.79 22.41/12.35 21.62/15.54 19.48/12.88 17.42/12.23
Query 17.02/13.52 17.77/12.49 17.61/12.62 17.59/12.93 19.33/14.71 20.77/14.38 20.72/12.91 18.35/13.33
Combined 16.38/12.93 17.56/12.12 17.01/12.69 14.82/10.73 16.70/10.33 20.97/15.00 25.52/14.93 17.95/12.66
Phi-4-Multimodal Audio 21.86/15.08 20.72/14.57 14.56/10.01 18.17/13.53 18.75/11.38 22.95/16.53 18.72/10.68 20.04/13.72
Query 24.38/18.57 17.01/13.37 16.69/11.45 12.82/08.58 15.82/10.70 29.85/17.40 17.31/10.52 19.93/13.95
Combined 19.94/12.96 16.91/11.63 17.91/12.59 18.23/12.59 15.80/09.79 20.68/14.80 22.31/15.72 18.79/12.76
Retrieval Models
CLAP Audio 22.38/15.68 21.22/15.14 18.86/14.93 15.12/12.11 26.42/16.25 20.56/15.41 22.51/13.40 21.15/14.90
Query 23.56/16.87 23.45/16.67 18.50/13.78 18.23/12.82 23.74/15.78 23.43/17.09 26.56/15.98 22.71/15.90
Combined 23.79/16.96 22.10/15.85 18.52/15.24 17.26/12.19 25.00/16.43 22.14/17.18 27.61/15.23 22.43/15.82
CoLLAP Audio 17.23/12.22 19.90/14.49 19.75/13.52 19.04/12.75 20.24/14.11 18.17/14.13 27.75/15.52 19.62/13.62
Query 19.26/12.70 20.30/14.75 20.38/15.96 22.11/13.87 20.19/12.07 21.65/16.44 25.14/14.35 20.85/14.14
Combined 18.68/12.86 21.09/15.08 21.88/15.98 21.17/14.63 22.35/13.04 18.42/12.66 23.08/14.02 20.52/13.92
Traditional Methods
Popularity Query 16.41/11.21 15.15/10.18 16.52/12.30 16.73/10.10 14.92/9.62 15.05/11.67 23.59/13.70 16.51/11.09
Neighborhood Query 20.12/13.53 17.47/10.92 15.85/10.35 13.76/7.92 11.91/6.67 8.96/5.59 4.30/2.38 14.72/9.30
Table 1: Performance comparison across music genres and input modalities. Values represent Recall@20/nDCG@20. Bold values indicate the best-performing modality configuration for each model, and underlined values show the top-performing model for each genre, both determined by Recall@20.
3.2 Research Questions and Analysis
Our comprehensive evaluation addresses three key research questions that illuminate fundamental challenges in multimodal music recommendation (Table 1):
RQ1: How do different input modalities impact music recommendation performance?
Our analysis reveals that multimodal approaches do not consistently outperform single-modality configurations. CLAP achieves optimal performance with query-only input (22.7122.71% Recall@20) compared to audio-only (21.1521.15%), while Qwen2.5-Omni demonstrates superior performance with audio+query combination (21.9321.93% vs. 19.2619.26% audio-only and 18.2418.24% query-only). However, only 2 of 9 multimodal-capable models achieve their best performance in multimodal settings. This finding aligns with recent work [26, 27] demonstrating that current multimodal models exhibit architectural limitations in cross-modal integration, suggesting that existing fusion mechanisms inadequately exploit complementary information across modalities. A representative example is provided in Box 2.2.
Fig. 4: Mean Reciprocal Rank (MRR) comparison across genres and model types.
RQ2: How do different model paradigms compare in music recommendation?
On average, retrieval-based approaches demonstrate superior performance compared to generative models across all metrics (Recall@20: 21.2121.21% vs. 18.7718.77%; nDCG@20: 14.7214.72% vs. 13.1413.14%). CLAP achieves the highest overall performance (22.7122.71% Recall@20), while the best-performing generative model (Qwen2.5-Omni) reaches a competitive 21.9321.93%, surpassing CoLLAP (20.8520.85%). Traditional methods exhibit substantially lower performance, with popularity-based recommendation achieving 16.5116.51% and neighbourhood-based recommendation 14.7214.72%, highlighting their inadequacy for audio-centric recommendation tasks that require understanding musical content beyond popularity signals. Within generative models, performance varies from 15.7715.77% to 21.9321.93% across modality configurations, whereas retrieval models range from 19.6219.62% to 22.7122.71%. These results indicate that learned embedding representations provide more robust recommendation signals than generative reasoning for music recommendation tasks; see Case Study in Box 3.
RQ3: What genre-specific patterns emerge?
Performance analysis (Fig. 4) reveals substantial genre-specific variation reflecting both domain-specific challenges and model capabilities. Jazz consistently achieves strong performance across most models (28.0928.09% Recall@20 for Qwen2.5-Omni audio-only; CLAP: 27.6127.61% combined, CoLLAP: 27.7527.75%), while Pop exhibits the lowest peak performance across genres (23.3823.38%). Metal shows notable model-dependent variation, with CLAP achieving 26.4226.42% in audio-only configuration while other models show more moderate results (e.g., Qwen2.5-Omni combined: 23.6123.61%). Classical demonstrates a distinctive pattern where multimodal approaches substantially outperform single-modality inputs (Qwen2.5-Omni: 26.5326.53% combined vs. 14.8014.80% audio-only and 14.0714.07% query-only), suggesting that cross-modal integration may be particularly beneficial for this genre. These patterns indicate that genre-specific audio characteristics and linguistic conventions differentially impact model performance, with Jazz’s distinctive musical features being more readily captured by current architectures.
4 Conclusion
We introduced MusiCRS, the first benchmark for audio-centric conversational recommendation grounded in both conversational context and corresponding audio. Extensive evaluations across recommendation paradigms and genres reveal key limitations of current multimodal and conversational models, particularly in nuanced audio-contextual reasoning. We release MusiCRS to support future advances in multimodal music recommendation.
5 Acknowledgment
This work was partially supported by the U.S. National Science Foundation under Grant IIS-2432486.